OpenStudy (anonymous):

Find w(7) for the function w(x) = 5x + 2. 59 45♥ 37 13 Check my work Please :] "♥" marks what i chose

OpenStudy (oaktree):

Not quite. Look:

OpenStudy (oaktree):

w(7) = 5(7)+2 = 35+2 = ?

OpenStudy (anonymous):

Ohhhh so 37

OpenStudy (oaktree):

Yup!